please help me to solve this equation------------------ How much must be placed in a five-month term deposit earning 4.5% simple interest in order to earn $510 interest? (Round your answer to the nearest cent.)
Answer by stanbon(75887)

How much must be placed in a five-month term deposit earning 4.5% simple interest in order to earn $510 interest? (Round your answer to the nearest cent.)
-----
I = P*r*t
-----
510 = P*0.045*(5/12)
----
P = 510/(0.045(5/12))
---
P = 510/0.01875
----
P = $27,200.00
----------
Cheers,
Stan H.
-------------